# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 4 Numerical Results
In this example, we consider a problem in $\Omega=[-5,5]$ to verify our convergence analysis of the fictitious play, specifically Lemma 2.10 and the linear convergence. We take non-zero viscosity $\nu=0.1$, quadratic Hamiltonian $H(x,p)=\frac{1}{2}\|p\|_{2}^{2}$, local interaction cost $f(x,\rho)=\rho(x)$, fixed terminal cost $f_{T}(x,\rho)=0$ and choose the initial density $\rho_{0}(x)=\rho_{G}(x;0,0.5)$. For the discrete system, we take $n_{x}=10^{4}$ points in the space and $n_{t}=30$ in time. We choose the numerical viscosity to be $\nu_{n}=1$. The algorithm is initialized with $\phi^{(0)}=\mathbf{0}$ which implies a zero velocity field. In this example, we take the weight $\delta_{k}=0.5$ being a constant along iterations and set the stopping criteria $\epsilon=10^{-12}$.
@@
According to Lemma 2.10, $\langle\hat{\rho}-\rho^{*},\rho-\rho^{*}\rangle\leq 0$, achieving equality only when $\hat{\rho}=\rho^{*}$. Actually the plot of $\frac{\langle\rho^{(k)}-\rho^{*},\hat{\rho}^{(k)}-\rho^{*}\rangle_{\mathcal{G}}}{\|\rho^{(k)}-\rho^{*}\|_{\mathcal{G}}\|\hat{\rho}^{(k)}-\rho^{*}\|_{\mathcal{G}}}$ in Figure 3 demonstrates that $\hat{\rho}^{(k)}-\rho^{*}$ and $\rho^{(k)}-\rho^{*}$ are almost in perfect opposite directions which makes fictitious play stable and efficient when using an appropriate convex weighted average of $\hat{\rho}^{(k)}$ and $\rho^{(k)}$ for an update. Additionally, the comparison between $\delta_{k}=0.5$ and $\delta_{k}=1$ suggests that more aggressive updates may not be a good choice. In particular, $\delta_{k}=1$ corresponds to the fixed point iteration which generates a sequence that is jumping left and right without convergence for this particular example.
@@ 4.2 Convergence rate independent of mesh step
This example aims to demonstrate that the convergence of fictitious play is independent of the mesh size and the discretization accuracy which is a motivation for using hierarchical grids and the assumption for the complexity analysis in table 1. On the other hand, since the constraint involves differential operators, if an optimization method is used to solve a potential MFG, the condition number of the discretized system becomes larger and larger when the mesh size becomes smaller and smaller, which causes slower and slower convergence.
We consider a non-potential MFG on $\Omega=[-5,5]^{2}$ with standard Hamiltonian $H(x,p)=\frac{1}{2}|p|^{2}$ and viscosity $\nu=1$. The interaction cost is a convolution $f(x,\rho)=10\int_{\Omega}K(x-y)\rho(y)\mathrm{d}y$, where the kernel is a non-isometric Gaussian $K(x)=\rho_{G}(x_{1};0,4)\rho_{G}(x_{2};0,0.5)$ as illustrated in Figure 4. Due to this interaction cost, the population tends to avoid concentration along the $x_{1}$ direction. We choose the initial and desired terminal density to be $\rho_{0}(x)=\rho_{G}(x_{1};-2,1)\rho_{G}(x_{2};-2,0.5)$, $\rho_{1}(x)=\rho_{G}(x_{1};2,1)\rho_{G}(x_{2};2,0.5)$. As illustrated in Figure 4, they are non-isometric Gaussians as well. Although the initial and terminal distributions have a larger variance along the $x_{1}$ direction, the interaction kernel induces a larger cost along the $x_{1}$ direction. To reduce the interaction cost, the players tend to avoid aggregating along the $x_{1}$ direction. The terminal cost is $f_{T}(x,\rho)=150(\rho-\rho_{1})$ which pushes the terminal density to the desired one.
@@ 4.3 Acceleration with back-tracking line-search
We discretize the space with $n_{x}=2\times 10^{3}$ segments, the time domain with $n_{t}=100$ segments, and take numerical viscosity $\nu_{n}=0$. The algorithm initializes with zero velocity field and the tolerance is $\epsilon=10^{-10}$. We select $\delta_{k}$ by back-tracking line-search method in Section 2.2.4 and the parameters are $\delta=1,\beta=0.5$ and $\zeta=0.8$. The algorithm with line-search weight converges in 133 iterations in 23.13s. We also run the algorithm with diminishing weight $\delta_{k}=\frac{2}{k+2}$ and constant weight $\delta_{k}=0.1$, $\delta_{k}=0.5$ for 180 iterations and compare the convergence in Figure 7, where “dimi” stands for diminishing weight and “btls” stands for back-tracking line-search weight. Figure 7 shows that back-tracking line-search efficiently improves the convergence compared to diminishing weight or non-optimal pre-select constant weight (0.1), and it also avoids getting stuck at some point with an aggressive pre-select constant weight (0.5). We remark that the oscillation in the consecutive residue and Fokker-Planck residue is reasonable since the back-tracking line search only guarantees the decay of $g^{(k)}$ and how the consecutive residue is bounded by $g^{(k)}$ depends on the local convexity property between $\rho^{(k-1)}$ and $\hat{\rho}^{(k)}$, which can be different for different $k$. Since the weight sequence generated by the back-tracking line search is not smooth, the oscillations occur. This oscillation is not observed for diminishing weight or constant weight because the weight sequences are smoother there and therefore generate a smoother sequence of $\rho^{(k)}$.
@@
Let $\mu_{1}(\rho),\mu_{2}(\rho)$ be first and second-order moments of the distribution $\rho$. We consider a MFG on $\Omega=\mathbb{R}$ with initial distribution density $\rho_{0}(x)=\rho_{G}(x;c,\sigma_{0})$, Hamiltonian $H(x,p)=\frac{1}{2}|p|_{2}^{2}$, interaction cost $f(x,t,\rho)=\frac{1}{2}\left(\alpha^{2}+\frac{\nu^{2}}{(\mu_{2}(\rho)-\mu_{1}^{2}(\rho))^{2}}\right)\left(x-\mu_{1}(\rho)\right)^{2}+2\alpha x+\left(\frac{1}{2}(2at+b)^{2}+\alpha\nu-\frac{\nu^{2}}{\mu_{2}(\rho)-\mu_{1}^{2}(\rho)}\right)$, terminal cost $f_{T}(x,\rho)=-(2a+b)x-\frac{1}{2}\left(\alpha-\frac{\nu}{\mu_{2}(\rho)-\mu_{1}^{2}(\rho)}\right)\left(x-\mu_{1}(\rho)\right)^{2}$. Here $a,b,c,\sigma_{0},\alpha$ are pre-selected constants. Let $\mu(t)=at^{2}+bt+c$ and $\sigma(t)=\sigma_{0}\exp{(\alpha t)}$. A solution to this system is
@@
The first row of Figure 8 shows the gain function, the consecutive residue, and the FP residue against the accumulated number of iterations. The jumps are caused by the grid changes when the solution on a coarse grid is interpolated to a finer grid which will generate a larger error initially due to a change of discretization on different mesh sizes. Since the exact solution is available in this example, we plot the error against the accumulated number of iterations in the second row. It shows that the hierarchical grid fictitious play effectively converges to the Nash Equilibrium.
@@ 4.6 Mean-field planning with hierarchical grid strategy
In this example we solve a mean-field planning (MFP) problem (96) in $\Omega=[-5,5]^{2}\subset\mathbb{R}^{2}$. Different from (2), both boundary conditions of a MFP are on population density.
@@
As discussed in [1], one can solve the MFP (96) by considering the penalized MFG (97) with large positive $\eta$.
@@
Let the viscosity $\nu=1$. We would like to solve the problem with interaction cost $f(x,\rho(t))=\begin{cases}+\infty,&\|x\|_{2}^{2}\leq 2,\\ 0,&\text{otherwise}\end{cases}$ which is the indicator function of a disk and is independent of the population density. And we take the initial and terminal density to be two Gaussian distributions with means $[-3,-3]$ and $[3,3]$, respectively, and both with covariance matrix $\operatorname{diag}(0.25,0.25)$. The solution $\rho$ should avoid the disk and transport from one Gaussian distribution to the other.
We solve (97) on a hierarchy of grids with increasing penalty parameters. On the $l$-th level grid $\mathcal{G}_{l}$, the domain is discretized with $n_{t}=2\times 2^{l},n_{x_{1}}=n_{x_{2}}=8\times 2^{l}$. To simulate the indicator function and $\eta\to\infty$, on the $l$-th grid, we take $\eta=50l$ and solve (97) with $f(x,\rho(t))=\begin{cases}2\eta,&\|x\|_{2}^{2}\leq 2,\\ 0,&\text{otherwise}.\end{cases}$ At each level, we run Algorithm 3 until tolerance $\epsilon=1\times 10^{-6}$ is achieved or maximal 200 iteration is achieved. The initialization of the $l$-th layer is the prolongation of the output on the $l-1$-th layer.
@@ 5 Conclusion
This paper studies the convergence and acceleration of fictitious play. We show that under mild assumptions, fictitious play converges linearly. The convergence analysis does not rely on the existence of a potential and is therefore applicable for general non-potential MFGs. Our convergence analysis shows that the weight is crucial to the convergence rate and therefore we propose a back-tracking line search strategy to choose the best weight in every iteration. In addition, we speed up the algorithm by working with a hierarchy of grids to accelerate the information transfer in the time and space joint domain. We are exploring the efficient implementation of fictitious play in high-dimension space and the convergence analysis when the intermediate optimization cannot be solved exactly.
@@ Acknowledgement
J.Y. and X.C. were partially supported by NSF DMS-2237842 and Simons Foundation (grant ID: 814643). J-G.L. was supported by NSF DMS-2106988. H. Zhao was partially supported by NSF DMS-2012860 and DMS-2309551.

### v2

@@ 5.1 Divergence of fixed-point iteration
In this example, we consider a problem on $[-5,5]$ to verify our convergence analysis of the fictitious play, specifically Proposition 3.11 and the linear convergence. We take non-zero viscosity $\nu=0.1$, quadratic Hamiltonian $H(x,p)=\frac{1}{2}|p|_{2}^{2}$, local interaction cost $f(x,\rho)=\rho(x)$, fixed terminal cost $f_{T}(x,\rho)=0$ and choose the initial density $\rho_{0}(x)=\rho_{G}(x;0,0.5)$. For the discrete system, we take $n_{x}=10^{4}$ points in the space and $n_{t}=30$ in time. We choose the numerical viscosity to be $\nu_{n}=1$. The algorithm is initialized with $\phi^{(0)}=\mathbf{0}$ which implies a zero velocity field. In this example, we take the weight $\delta_{k}=0.5$ being a constant along iterations and set the stopping criteria $\epsilon=10^{-12}$.
@@
According to Proposition 3.11, $\langle\hat{\rho}-\rho^{*},\rho-\rho^{*}\rangle\leq 0$, achieving equality only when $\hat{\rho}=\rho^{*}$. Actually the plot of $\frac{\langle\rho^{(k)}-\rho^{*},\hat{\rho}^{(k)}-\rho^{*}\rangle_{\mathcal{G}}}{\|\rho^{(k)}-\rho^{*}\|_{\mathcal{G}}\|\hat{\rho}^{(k)}-\rho^{*}\|_{\mathcal{G}}}$ in Figure 4 demonstrates that $\hat{\rho}^{(k)}-\rho^{*}$ and $\rho^{(k)}-\rho^{*}$ are almost in perfect opposite directions which makes fictitious play stable and efficient when using an appropriate convex weighted average of $\hat{\rho}^{(k)}$ and $\rho^{(k)}$ for an update. Additionally, the comparison between $\delta_{k}=0.5$ and $\delta_{k}=1$ suggests that more aggressive updates may not be a good choice. In particular, $\delta_{k}=1$ corresponds to the fixed point iteration, which generates a sequence that jumps left and right without converging for this particular example.
@@ 5.3 Convergence rate independent of mesh step
This example aims to demonstrate that the convergence of fictitious play is independent of the mesh size and the discretization accuracy, which is a motivation for using hierarchical grids and the assumption for the complexity analysis in Table 2. On the other hand, since the constraint involves differential operators, if an optimization method is used to solve a potential MFG, the condition number of the discretized system becomes larger when the mesh size becomes smaller, which causes slower convergence.
We consider a non-potential MFG on $[-5,5]^{2}$ with standard Hamiltonian $H(x,p)=\frac{1}{2}|p|^{2}$ and viscosity $\nu=1$. The interaction cost is a convolution $f(x,\rho)=10\int_{\mathbb{T}^{d}}K(x-y)\rho(y)\mathrm{d}y$, where the kernel is a non-isometric Gaussian $K(x)=\rho_{G}(x_{1};0,4)\rho_{G}(x_{2};0,0.5)$ as illustrated in Figure 5. Due to this interaction cost, the population tends to avoid concentration along the $x_{1}$ direction. We choose the initial and desired terminal density to be $\rho_{0}(x)=\rho_{G}(x_{1};-2,1)\rho_{G}(x_{2};-2,0.5)$, $\rho_{1}(x)=\rho_{G}(x_{1};2,1)\rho_{G}(x_{2};2,0.5)$. As illustrated in Figure 5, they are non-isometric Gaussians as well. Although the initial and terminal distributions have a larger variance along the $x_{1}$ direction, the interaction kernel induces a larger cost along the $x_{1}$ direction. To reduce the interaction cost, the players tend to avoid aggregating along the $x_{1}$ direction. The terminal cost is $f_{T}(x,\rho)=150(\rho-\rho_{1})$, which pushes the terminal density to the desired one.
@@ 5.4 Acceleration with back-tracking line-search
We discretize the space with $n_{x}=2\times 10^{3}$ segments, the time domain with $n_{t}=100$ segments, and take numerical viscosity $\nu_{n}=0$. The algorithm initializes with a zero velocity field and the tolerance is $\epsilon=10^{-10}$. We select $\delta_{k}$ by the back-tracking line-search method in Section 3.5 and the parameters are $\delta=1,\beta=0.5$ and $\zeta=0.8$. The algorithm with line-search weight converges in 133 iterations in 23.13s. We also run the algorithm with diminishing weight $\delta_{k}=\frac{2}{k+2}$ and constant weight $\delta_{k}=0.1$, $\delta_{k}=0.5$ for 180 iterations and compare the convergence in Figure 8, where “dimi” stands for diminishing weight and “btls” stands for back-tracking line-search weight. Figure 8 shows that back-tracking line-search efficiently improves the convergence compared to diminishing weight or non-optimal pre-select constant weight (0.1), and it also avoids getting stuck at some point with an aggressive pre-select constant weight (0.5). We remark that the oscillation in the consecutive residue and Fokker-Planck residue is reasonable since the back-tracking line search only guarantees the decay of $g^{(k)}$ and how the consecutive residue is bounded by $g^{(k)}$ depends on the local convexity property between $\rho^{(k-1)}$ and $\hat{\rho}^{(k)}$, which can be different for different $k$. Since the weight sequence generated by the back-tracking line search is not smooth, oscillations occur. This oscillation is not observed for diminishing weight or constant weight because the weight sequences are smoother there and therefore generate a smoother sequence of $\rho^{(k)}$.
@@
Let $\mu_{1}(\rho),\mu_{2}(\rho)$ be first and second-order moments of the distribution $\rho$. We consider a MFG on $\mathbb{R}$ with initial distribution density $\rho_{0}(x)=\rho_{G}(x;c,\sigma_{0})$, Hamiltonian $H(x,p)=\frac{1}{2}|p|_{2}^{2}$, interaction cost $f(x,t,\rho)=\frac{1}{2}\left(\alpha^{2}+\frac{\nu^{2}}{(\mu_{2}(\rho)-\mu_{1}^{2}(\rho))^{2}}\right)\left(x-\mu_{1}(\rho)\right)^{2}+2\alpha x+\left(\frac{1}{2}(2at+b)^{2}+\alpha\nu-\frac{\nu^{2}}{\mu_{2}(\rho)-\mu_{1}^{2}(\rho)}\right)$, terminal cost $f_{T}(x,\rho)=-(2a+b)x-\frac{1}{2}\left(\alpha-\frac{\nu}{\mu_{2}(\rho)-\mu_{1}^{2}(\rho)}\right)\left(x-\mu_{1}(\rho)\right)^{2}$. Here $a,b,c,\sigma_{0},\alpha$ are pre-selected constants. Let $\mu(t)=at^{2}+bt+c$ and $\sigma(t)=\sigma_{0}\exp{(\alpha t)}$. A solution to this system is
@@
The first row of Figure 9 shows the gain function, the consecutive residue, and the FP residue against the accumulated number of iterations. The jumps are caused by the grid changes when the solution on a coarse grid is interpolated to a finer grid, which will generate a larger error initially due to a change of discretization on different mesh sizes. Since the exact solution is available in this example, we plot the error against the accumulated number of iterations in the second row. It shows that the hierarchical grid fictitious play effectively converges to the Nash Equilibrium.
@@ 5.7 Mean-field planning with hierarchical grid strategy
In this example, we solve a mean-field planning (MFP) problem (108) in $[-5,5]^{2}\subset\mathbb{R}^{2}$. Different from (2), both boundary conditions of a MFP are on population density.
@@
As discussed in Achdou2012mfp , one can solve the MFP (108) by considering the penalized MFG (109) with large positive $\eta$.
@@
Let the viscosity $\nu=1$. We would like to solve the problem with interaction cost $f(x,\rho(t))=\begin{cases}+\infty,&|x|_{2}^{2}\leq 2,\\ 0,&\text{otherwise}\end{cases}$, which is the indicator function of a disk and is independent of the population density. And we take the initial and terminal density to be two Gaussian distributions with means $[-3,-3]$ and $[3,3]$, respectively, and both with covariance matrix $\operatorname{diag}(0.25,0.25)$. The solution $\rho$ should avoid the disk and transport from one Gaussian distribution to the other.
We solve (109) on a hierarchy of grids with increasing penalty parameters. On the $l$-th level grid $\mathcal{G}_{l}$, the domain is discretized with $n_{t}=2\times 2^{l},n_{x_{1}}=n_{x_{2}}=8\times 2^{l}$. To simulate the indicator function and $\eta\to\infty$, on the $l$-th grid, we take $\eta=50l$ and solve (109) with $f(x,\rho(t))=\begin{cases}2\eta,&|x|_{2}^{2}\leq 2,\\ 0,&\text{otherwise}.\end{cases}$ At each level, we run Algorithm 3 until tolerance $\epsilon=1\times 10^{-6}$ is achieved or a maximum of 200 iterations is achieved. The initialization of the $l$-th layer is the prolongation of the output on the $l-1$-th layer.
@@ 6 Conclusion
This paper studies the convergence and acceleration of fictitious play. We show that under mild assumptions, fictitious play converges linearly. The convergence analysis does not rely on the existence of a potential and is therefore applicable to general non-potential MFGs. Our convergence analysis shows that the weight is crucial to the convergence rate, and therefore, we propose a back-tracking line search strategy to choose the best weight in every iteration. In addition, we speed up the algorithm by working with a hierarchy of grids to accelerate the information transfer in the time and space joint domain. We are exploring the efficient implementation of fictitious play in high-dimensional space and the convergence analysis when the intermediate optimization cannot be solved exactly.
@@ Acknowledgement
J.Y. and X.C. were partially supported by NSF DMS-2237842 and Simons Foundation (grant ID: MPS-MODL-00814643). J-G.L was supported by NSF DMS-2106988. H. Zhao was partially supported by NSF DMS-2012860 and DMS-2309551.